How many electrons constitute 1C of charge..............?

Grade:12

6 Answers

SAI SARDAR
8 years ago
We know that the a electron have charge of 1.6 × 10−19 C. then  6.65 * 1018 electrons constitute  a coulomb of charge.
SHANMUKESHWAR
8 years ago
e =1.6 × 10−19 C. Total charge required 1Coulomb.
the number of electrons constituting one coulomb of charge. We know that the a electron have charge of 1.6 × 10−19 C.
Karthik Kumar
8 years ago
1 electron have charge of 1.6 *10-19 C
1 coulomb of charge contains=1/(1.6 *10-19) =6.24*1018 electrons
Soni Mishra
7 years ago
in 1.6*10^-19C charge there is one electron 
so,
in 1c there will be 1/(1.6*10^-19)electrons=6.25*10^18 electrons
sandip ghimire
5 years ago
we know that the electron have charge of 1.6*10^-19C 
then 1 coulomb of charge contans = 1/(1.6*10^-19)=6.25*10^18 electrons.
 
Yash Chourasiya
3 years ago
Dear Student

We Know that
e = 1.6 × 10-19 C
Total charge required 1Coulomb.
Therefore q = 1C
We know the formula

q=ne
n =q/e—–(i)

Hence, 1 C charge is carried by

Substituting the values in equation (i) we get,

1/ (-1.6⨯10-19) = 6.25 ⨯ 1018 electrons.

I hope this answer will help you.
Thanks & Regards
Yash Chourasiya